OpenStudy (anonymous):

WILL MEDAL AND FAN BEST ANSWER!! i need explanation AND answer. Charmaine works mowing lawns and babysitting. She earns $8.20 an hour for mowing and $7.70 an hour for babysitting. How much will she earn for 6 hours of mowing and 1 hour of babysitting?

OpenStudy (jennyrlz):

hello :)

OpenStudy (jennyrlz):

lets turn the problem into a much simpler looking equation :)

OpenStudy (jennyrlz):

8.20x+7.70y=Answer

OpenStudy (jennyrlz):

Are you with me?

OpenStudy (anonymous):

no, its hard with letters

OpenStudy (jennyrlz):

ok well i turned it into a simple equation,. Ijust want ed to make sure you were not "AFK"

OpenStudy (jennyrlz):

X and Y stand for the hours

OpenStudy (jennyrlz):

8.20(6)+7.70(1)= answer

OpenStudy (jennyrlz):

do you see where i am getting these numbers from?

OpenStudy (jennyrlz):

Well next you just "solve" it

OpenStudy (jennyrlz):

8.20 x 6 = 49.2 49.2 + 7.7 = answer

OpenStudy (jennyrlz):

the answer is 56.9 any questions message me when you get back

OpenStudy (anonymous):

ty